# OpenMathInstruct-1: A 1.8 Million Math Instruction Tuning Dataset

### v1

@@ Appendix A Data Statistics
A.1 MATH Training Data Split by Subjects
@@
A.2 Code-Block Count Frequencies
Let $d$ and $e$ denote the solutions of $02x^{2}+3x-5=$ 0. What is the value of $(d-1)(e-1)$?
Let’s solve the quadratic equation using Sympy: <llm-code> from sympy import Symbol, solve, Eq, simplify # Define the variable x x = Symbol(’x’) # Define the equation eq = 2*x**2 + 3*x - 5 # Solve the equation roots = solve(eq, x) # Print the solutions print("The solutions are:") for root in roots:     print("x = ", simplify(root)) </llm-code> <llm-code-output> The solutions are: x =  -5/2 x =  1 </llm-code-output> So $d=-5/2$ and $e=1$. Let’s calculate the answer: <llm-code> d = -5/2 e = 1 d_minus_1 = d - 1 e_minus_1 = e - 1 result = d_minus_1 * e_minus_1 print("The value of (d-1)(e-1) is:", result) </llm-code> <llm-code-output> The value of (d-1)(e-1) is: -0.0 </llm-code-output> So the answer is $\boxed{-0.0}$.
Let $p(x)=\sqrt{-x}$, and $q(x)=8x^{2}+10x-3$. The domain of $p(q(x))$ can be written in the form $a\leq x\leq b$. Find $b-a$.
Lets use sympy to solve for the domain of \$p(q(x))\$. <llm-code> import sympy as sp # define the symbols x = sp.symbols(’x’) # define the functions p = sp.sqrt(-x) q = 8*x**2 + 10*x - 3 # solve for the domain of p(q(x)) domain = sp.solve(q >= 0, x) # print the domain print(domain) </llm-code> <llm-code-output> ((1/4 <= x) & (x < oo)) | ((-oo < x) & (x <= -3/2)) </llm-code-output> So the domain is \$x $\in$ [1/4, $\infty$) $\cup$ ($\infty$, -3/2)\$. The difference between the upper and lower bounds is $b-a=\boxed{1/4-(-3/2)}$.
Their 1st child was born 15 years ago, and therefore is 15 years old. Their 2nd child was born 1 year after their 15-year-old child, and therefore is 15-1=14 years old. Their 3rd child was born 4 years after their 14-year-old child, and therefore is 14-4=10 years old. Their 4th child was born 2 years after their 10-year-old child, and therefore is 10-2=8 years old. Let’s write down a python script to answer this problem. <llm-code> children = [’first’, second’, third’, fourth’] child_age = [15, 14, 10, 8] number_of_children = len(children) children = children[:-1] # lets get rid of the youngest one since we already know that one dictionary = dict(zip(children, child_age)) dictionary </llm-code> <llm-code-output> {’first’: 15, second’: 14, third’: 10} </llm-code-output> The answer is 888 years old.
Let $f(x)$ be the number of lassis she can make out of $x$ mangoes. From the question, we can see that $f(2)=11$. Using basic algebra, we can see that $f(12)=\boxed{66}$. Lets verify this with sympy. <llm-code> import sympy as sp # define the unknown function x, y = sp.symbols(’x y’) # lets define the parabola parabola = sp.Eq(y, x**2 + b*x + c) # substitute points into parabola equation and solve for b, c point_1 = parabola.subs({x: -1, y: -11}) point_2 = parabola.subs({x: 3, y: 17}) solutions = sp.solve((point_1,point_2), (b, c)) solutions[b] </llm-code>
Let $f(x)$ be the number of lassis she can make out of $x$ mangoes. From the question, we can see that $f(2)=11$. Using basic algebra, we can see that $f(12)=\boxed{66}$.
The areas of two squares are in the ratio $25:36$. What is the ratio of their perimeters? Express your answer in the form $a:b$.
In general, if the ratio of the perimeters of two similar figures is $a/b$, then the ratio of their areas is $(a/b)^{2}$. So in this case, $(a/b)^{2}=25/36$, or $a/b=5/6$. So the answer is $\boxed{5:6}$. Let’s use sympy to print out the difference between the ratio of their perimeters and each of the options. <llm-code> from sympy import Rational, Abs # areas are in the ratio 25:36 area_ratio = Rational(25, 36) # list of options options = [Rational(5, 6), Rational(5, 4), Rational(5, 3), Rational(5, 2), Rational(5, 1)] # lets print out the differences [Abs(area_ratio - frac_option) for frac_option in options] </llm-code> <llm-code-output> [5/36, 5/9, 35/36, 65/36, 155/36] </llm-code-output> Let’s now check which difference is the smallest. <llm-code> import numpy as np # Calculate the idx of the closest option min_idx = np.argmin([5/36, 5/9, 35/36, 65/36, 155/36]) # Print the closest option print(options[min_idx]) </llm-code> <llm-code-output> 5/6 </llm-code-output> So the answer is 5:6:565 : 6.
The code-interpreter allows solving a problem by breaking it into multiple code blocks. Most of the solutions in OpenMathInstruct-1 have 0 or 1 code blocks, 16.4% and 81.7% of the solutions, respectively. The remaining 2% of the solutions have two or more code blocks. Figure 7 shows an example of a solution using two code blocks.
(a) GSM8K
(b) MATH
Figure 12: Histogram of the number of solutions for problems in GSM8K and MATH.
A.3 Frequency of solutions per problem in OpenMathInstruct-1
Figure 12 shows the histogram of the frequency of the number of solutions in OpenMathInstruct-1 for problems in GSM8K and MATH. For GSM8K, 57.4% problems (4292 out of 7473) have more than half the sampled solutions, i.e., $>$ 128 solutions out of 256 sampled solutions, that are valid. For MATH, the solution frequency histogram mimics a power-law curve, with most problems having very few solutions and very few problems with many solutions. Concretely, 19% problems (1324 out of 6978) with any solutions in OpenMathInstruct-1 have $\leq 10$ valid solutions out of the 896 sampled solutions per problem. To further highlight the hardness of the MATH benchmark in comparison to GSM8K, just 3.1% problems (233 out of 7500) have more than half the sampled solutions ( $>448$ out of 896 sampled solutions), that are valid.
Let’s use sympy to solve this problem. <llm-code> from sympy import symbols, Eq, solve # define the variables base, altitude = symbols(’base altitude’) # area of the triangle area_eq = Eq(base * altitude / 2, 600) # length of the corresponding base base_eq = Eq(base, 30) # solve the equations solutions = solve((area_eq, base_eq), (base, altitude)) # print the altitude print(solutions[altitude]) </llm-code> <llm-code-output> [0;31m————————–[0m [0;31mTypeError[0m                                 Traceback (most recent call last) File [0;32m<ipython-input-1-f95732badac7>:16[0m [1;32m     13[0m solutions [38;5;241m=[39m solve((area_eq, base_eq), (base, altitude)) [1;32m     15[0m [38;5;66;03m# print the altitude[39;00m [0;32m—> 16[0m [38;5;28mprint[39m([43msolutions[49m[43m[[49m[43maltitude[49m[43m][49m) [0;31mTypeError[0m: list indices must be integers or slices, not Symbol </llm-code-output> So the altitude is 202020.
What is the 100th digit to the right of the decimal point in the decimal representation of $\frac{13}{90}$?
[…]
We can choose a starting lineup in $12\times 11\times 10\times 9\times 8=\boxed{11880}$ ways.
@@ B.1 Training Hyperparameters
Table 11 details the hyperparameters used for finetuning the different base models. Due to compute limitations, for 70B models, we limit the training to 2 epochs for Llama2 and 3 epochs for CodeLlama. For CodeLlama models, we use the learning rate of 2e-5 used by ToRA models (Gou et al., 2024). For the Mistral-7B model, we use a lower learning rate of 1e-6 based on our hyperparameter search.
@@ B.2 Sample Solutions
  • Figure 11 shows a sample solution where the generated solution gets the right answer but through flawed reasoning. These semantically noisy solutions are much harder to detect with simple syntactic filters. One solution might be to use models like GPT-4 to grade the generated solutions as done in recent work (Gunasekar et al., 2023). We leave the work of developing such semantic filters for future work.
@@ B.3 Error Analysis of Solutions Generated by Fine-tuned Model
In this section, we illustrate instances of the different kind of errors made by the ablation model analyzed in Section 5.
@@
  • Figure 14 demonstrates a sample where the model performs correct reasoning while generating the code. However, the model falters at copying the code output and ends up generating a new answer.
  • Figure 15 illustrates a sample where the model performs correct reasoning but falters in arithmetic calculation (multiplication). Failure at arithmetic computation has been a known issue with LLMs and justifies our choice of preferring code-based solutions.
@@ B.4 Instructions for Few-shot Data Generation
Few-shot prompt text masking ($\mathcal{I}_{\text{mask}}$)
Here are some examples of questions, solutions, and their masked solutions followed by a new question and solution that you need to mask. The goal is to ensure that the masked solution doesn’t have any of the numerical values not mentioned in the question. So intermediate values calculated in the solution are to be masked by single letter capital variables, such as M, N.
Zero-shot prompt for fine-tuned model

### v2

@@ Appendix A Data Statistics
Sample Code-Interpreter Style Solution
Question A department store displays a 20% discount on all fixtures. What will be the new price of a 25 cm high bedside lamp that was worth \$120?
Code-Interpreter Style Solution Let’s solve this problem using Python code. <llm-code> discount_percent = 20 price_before_discount = 120 discount = discount_percent / 100 discount_amount = price_before_discount * discount price = price_before_discount - discount_amount price </llm-code> <llm-code-output> 96.0 </llm-code-output> So the new price of the lamp is 969696 dollars.
Solution with Two Code Blocks
Question Let $d$ and $e$ denote the solutions of $02x^{2}+3x-5=$ 0. What is the value of $(d-1)(e-1)$?
Generated Solution Let’s solve the quadratic equation using Sympy: <llm-code> from sympy import Symbol, solve, Eq, simplify # Define the variable x x = Symbol(’x’) # Define the equation eq = 2*x**2 + 3*x - 5 # Solve the equation roots = solve(eq, x) # Print the solutions print("The solutions are:") for root in roots: print("x = ", simplify(root)) </llm-code> <llm-code-output> The solutions are: x = -5/2 x = 1 </llm-code-output> So $d=-5/2$ and $e=1$. Let’s calculate the answer: <llm-code> d = -5/2 e = 1 d_minus_1 = d - 1 e_minus_1 = e - 1 result = d_minus_1 * e_minus_1 print("The value of (d-1)(e-1) is:", result) </llm-code> <llm-code-output> The value of (d-1)(e-1) is: -0.0 </llm-code-output> So the answer is $\boxed{-0.0}$.
A.1 Code-Block Count Frequencies
We’re using the code-interpreter solution format, which allows for flexible reasoning in text along with precision in code-based reasoning (see Figure 11). The code-interpreter allows solving a problem by breaking it into multiple code blocks. Most of the solutions in OpenMathInstruct-1 have 0 or 1 code blocks, 16.4% and 81.7% of the solutions, respectively (see Figure 7). The remaining 2% of the solutions have two or more code blocks. Figure 8 shows an example of a solution using two code blocks.
Diversity of Solutions
Question Let $t(x)=\sqrt{3x+1}$ and $f(x)=5-t(x)$. What is $t(f(5))$?
[…]
Figure 10: Three representative solutions for a sample problem from MATH training set in OpenMathInstruct-1. The first and third solutions use code, with the first one using sympy library. The second solution is text-based.
A.2 Frequency of solutions per problem in OpenMathInstruct-1
Figure LABEL:fig:hist shows the histogram of the frequency of the number of solutions in OpenMathInstruct-1 for problems in GSM8K and MATH. For GSM8K, 57.4% problems (4292 out of 7473) have more than half the sampled solutions, i.e., $>$ 128 solutions out of 256 sampled solutions, that are valid. For MATH, the solution frequency histogram mimics a power-law curve, with most problems having very few solutions and very few problems with many solutions. Concretely, 19% problems (1324 out of 6978) with any solutions in OpenMathInstruct-1 have $\leq 10$ valid solutions out of the 896 sampled solutions per problem. To further highlight the hardness of the MATH benchmark in comparison to GSM8K, just 3.1% problems (233 out of 7500) have more than half the sampled solutions ( $>448$ out of 896 sampled solutions), that are valid.
A.3 MATH Training Data Split by Subjects
@@ B.1 Training Hyperparameters
Table 12 details the hyperparameters used for finetuning the different base models. Due to compute limitations, for 70B models, we limit the training to 2 epochs for Llama2 and 3 epochs for CodeLlama. For CodeLlama models, we use the learning rate of 2e-5 used by ToRA models [8]. For the Mistral-7B model, we use a lower learning rate of 1e-6 based on our hyperparameter search.
Code + Text Reasoning
Question Let $p(x)=\sqrt{-x}$, and $q(x)=8x^{2}+10x-3$. The domain of $p(q(x))$ can be written in the form $a\leq x\leq b$. Find $b-a$.
Generated Solution Let’s use sympy to solve for the domain of $p(q(x))$. <llm-code> import sympy as sp # define the symbols x = sp.symbols(’x’) # define the functions p = sp.sqrt(-x) q = 8*x**2 + 10*x - 3 # solve for the domain of p(q(x)) domain = sp.solve(q >= 0, x) # print the domain print(domain) </llm-code> <llm-code-output> ((1/4 <= x) & (x < oo)) | ((-oo < x) & (x <= -3/2)) </llm-code-output> So the domain is $x\in[1/4,\infty)\cup(-\infty,-3/2)$. The difference between the upper and lower bounds is $b-a=\boxed{1/4-(-3/2)}$.
@@ B.2 Sample Solutions
  • Figure 10 presents a sample problem from the MATH training set with its representative solutions from the OpenMathInstruct-1 dataset.
@@
  • Figure 14 shows a sample solution where the generated solution gets the right answer but through flawed reasoning. These semantically noisy solutions are much harder to detect with simple syntactic filters. One solution might be to use models like GPT-4 to grade the generated solutions as done in recent work [5]. We leave the work of developing such semantic filters for future work.
@@ B.3 Error Analysis of Solutions Generated by Fine-tuned Model
In this section, we illustrate instances of the different kinds of errors made by the ablation model analyzed in Section 5.
  • Figure 15 demonstrates a sample where the model performs correct reasoning while generating the code. However, the model falters at copying the code output and generates a new answer.
  • Figure 16 presents an example where the model generates correct but inefficient code that doesn’t finish in our stipulated execution time and results in the timeout error.
  • Figure 17 illustrates a sample where the model performs correct reasoning but falters in arithmetic calculation (multiplication). Failure at arithmetic computation has been a known issue with LLMs and justifies our preference for code-based solutions.
  • Figure 18 shows an example where the model generates a wrong code block to solve the problem and repeatedly generates the same code block till it runs into our maximum execution constraint. This example shows a limitation of our current dataset where we lack examples in which the model recovers from its error, unlike GPT-4 [23].
@@
Shortcut Solution
Question Jolene and Phil have four children, each with the same birthday. They gave birth to their first child exactly 15 years ago. They gave birth to their second child exactly one year after the birth of their first child. They gave birth to their third child on the fourth birthday of their second child. Two years after the birth of their third child, they gave birth to their fourth child. How old, in years, is their fourth child?
Reference Solution Their 1st child was born 15 years ago, and therefore is 15 years old. Their 2nd child was born 1 year after their 15-year-old child, and therefore is 15-1=14 years old. Their 3rd child was born 4 years after their 14-year-old child, and therefore is 14-4=10 years old. Their 4th child was born 2 years after their 10-year-old child, and therefore is 10-2=8 years old.
Generated Solution Let’s write down a python script to answer this problem. <llm-code> children = [’first’, second’, third’, fourth’] child_age = [15, 14, 10, 8] number_of_children = len(children) children = children[:-1] # lets get rid of the youngest one since we already know that one dictionary = dict(zip(children, child_age)) dictionary </llm-code> <llm-code-output> {’first’: 15, second’: 14, third’: 10} </llm-code-output> The answer is 888 years old.
Solution Requiring Trimming
Question Caroline can make eleven lassis out of two mangoes. How many lassis can she make out of twelve mangoes?
Generated Solution Let $f(x)$ be the number of lassis she can make out of $x$ mangoes. From the question, we can see that $f(2)=11$. Using basic algebra, we can see that $f(12)=\boxed{66}$. Let’s verify this with sympy. <llm-code> import sympy as sp # define the unknown function x, y = sp.symbols(’x y’) # lets define the parabola parabola = sp.Eq(y, x**2 + b*x + c) # substitute points into parabola equation and solve for b, c point_1 = parabola.subs({x: -1, y: -11}) point_2 = parabola.subs({x: 3, y: 17}) solutions = sp.solve((point_1,point_2), (b, c)) solutions[b] </llm-code>
Trimmed Solution Let $f(x)$ be the number of lassis she can make out of $x$ mangoes. From the question, we can see that $f(2)=11$. Using basic algebra, we can see that $f(12)=\boxed{66}$.
Flawed Reasoning
Question The areas of two squares are in the ratio $25:36$. What is the ratio of their perimeters? Express your answer in the form $a:b$.
Reference Solution In general, if the ratio of the perimeters of two similar figures is $a/b$, then the ratio of their areas is $(a/b)^{2}$. So in this case, $(a/b)^{2}=25/36$, or $a/b=5/6$. So the answer is $\boxed{5:6}$.
Generated Solution Let’s use sympy to print out the difference between the ratio of their perimeters and each of the options. <llm-code> from sympy import Rational, Abs # areas are in the ratio 25:36 area_ratio = Rational(25, 36) # list of options options = [Rational(5, 6), Rational(5, 4), Rational(5, 3), Rational(5, 2), Rational(5, 1)] # lets print out the differences [Abs(area_ratio - frac_option) for frac_option in options] </llm-code> <llm-code-output> [5/36, 5/9, 35/36, 65/36, 155/36] </llm-code-output> Let’s now check which difference is the smallest. <llm-code> import numpy as np # Calculate the idx of the closest option min_idx = np.argmin([5/36, 5/9, 35/36, 65/36, 155/36]) # Print the closest option print(options[min_idx]) </llm-code> <llm-code-output> 5/6 </llm-code-output> So the answer is 5:6:565 : 6.
Last Step Mishap
Question What is the 100th digit to the right of the decimal point in the decimal representation of $\frac{13}{90}$?
[…]
Timeout
Question Let $a_{1},a_{2},\dots$ be a sequence for which $a_{1}=2$ , $a_{2}=3$, and $a_{n}=\frac{a_{n-1}}{a_{n-2}}$ for each positive integer $n\geq 3$. What is $a_{2006}$?
Generated Solution Let’s write a function that calculates $a_{n}$ for a given $n$. <llm-code> def a_n(n): if n == 1: return 2 elif n == 2: return 3 else: return a_n(n-1) / a_n(n-2) print(a_n(2006)) </llm-code> <llm-code-output> None </llm-code-output> So $a_{2006}=\boxed{1/5}$.
Figure 16: The model generates recursion-based code to solve the problem. While the code is technically correct, the lack of memoization means the code ends up being extremely inefficient and runs into timeout.
Calculation Error
Question Our basketball team has 12 members, each of whom can play any position. In how many ways can we choose a starting lineup consisting of a center, a power forward, a shooting forward, a point guard, and a shooting guard?
Generated Solution We can choose a starting lineup in $12\times 11\times 10\times 9\times 8=\boxed{11880}$ ways.
Maximum Executions Reached
Question In $\triangle ABC$, $AB=425$, $BC=450$, and $AC=510$. An interior point $P$ is then drawn, and segments are drawn through $P$ parallel to the sides of the triangle. If these three segments are of an equal length $d$, find $d$.
Generated Solution Let’s use sympy to solve this problem. <llm-code> from sympy import symbols, Eq, solve # Define the variables AB, BC, AC, d = symbols(’AB BC AC d’) # Define the equations eq1 = Eq(AB, 425) eq2 = Eq(BC, 450) eq3 = Eq(AC, 510) eq4 = Eq(AB + BC, 2 * d) eq5 = Eq(BC + AC, 2 * d) eq6 = Eq(AC + AB, 2 * d) # Solve the equations solutions = solve((eq1, eq2, eq3, eq4, eq5, eq6), (AB, BC, AC, d)) solutions </llm-code> <llm-code-output> [] </llm-code-output> The solutions are empty, so we can’t solve this problem using sympy. Let’s try to solve it manually. We can see that the sum of any two sides of a triangle is greater than the third side. So we can write the following inequalities: $AB+BC>AC$ $BC+AC>AB$ $AC+AB>BC$ Let’s rewrite them using the given values: $425+450>510$ $450+510>425$ $510+425>450$ We can solve these inequalities using sympy: $\cdots$
Figure 18: The model generates the first code block, which is logically incorrect and results in an empty output. The following text block identifies the problem with the output but performs incorrect reasoning leading up to the generation of the second code block. The model fails to recover from these errors and ends up generating code-text blocks that are verbatim copies of the first code-text block. The generation stops at the sixth code block, which is our limit on the number of code-block executions during inference.
Code Execution Error
Question The area of a triangle is 600 square feet. Find the altitude, in feet, of the triangle if the length of the corresponding base is 30 feet.
Generated Solution Let’s use sympy to solve this problem. <llm-code> from sympy import symbols, Eq, solve # define the variables base, altitude = symbols(’base altitude’) # area of the triangle area_eq = Eq(base * altitude / 2, 600) # length of the corresponding base base_eq = Eq(base, 30) # solve the equations solutions = solve((area_eq, base_eq), (base, altitude)) # print the altitude print(solutions[altitude]) </llm-code> <llm-code-output> [0;31m--------------------------[0m [0;31mTypeError[0m Traceback (most recent call last) File [0;32m<ipython-input-1-f95732badac7>:16[0m [1;32m 13[0m solutions [38;5;241m=[39m solve((area_eq, base_eq), (base, altitude)) [1;32m 15[0m [38;5;66;03m# print the altitude[39;00m [0;32m---> 16[0m [38;5;28mprint[39m([43msolutions[49m[43m[[49m[43maltitude[49m[43m][49m) [0;31mTypeError[0m: list indices must be integers or slices, not Symbol </llm-code-output> So the altitude is 202020.
@@ B.4 Instructions for Few-shot Data Generation
Few-shot prompt text masking ($\mathcal{I}_{\text{mask}}$) Here are some examples of questions, solutions, and their masked solutions followed by a new question and solution that you need to mask. The goal is to ensure that the masked solution doesn’t have any of the numerical values not mentioned in the question. So intermediate values calculated in the solution are to be masked by single letter capital variables, such as M, N.